Consider the following information:

Portfolio Expected Return Standard
Deviation
Risk-free 5.0 % 0 %
Market 10.6 23
A 8.6 12

a. Calculate the Sharpe ratios for the market portfolio and portfolio A. (Round your answers to 2 decimal places.)

Sharpe Ratio
Market portfolio
Portfolio A

b. If the simple CAPM is valid, is the above situation possible?

y/n
Answer #1
Answer a.
Calculation of Sharpe Ratio
Sharpe Ratio = (Rp - Rf) / SDp
Where,
Rp = Return of portfolio
Rf = Risk-free return
SDp = Standard deviation of portfolio
Now,
Sharpe ratio of Market Portfolio = (10.06 - 5) / 23
Sharpe ratio of Market Portfolio = 0.24
Share ratio of Portfolio A = (8.6 - 5) / 12
Sharpe Ratio of Portfolio A = 0.3
Answer b.
Since the Sharpe ratio is below 1, it means that the company has low expected return.
If CAPM is valid, then the company is paying stable returns as the company's required return is higher than the market premium.
Thus, the above situation is possible